The following reaction was monitored as a function of time:
AB→A+B
A plot of 1/[AB] versus time yields a straight line with slope 5.5×10−2 (M⋅s)−1 .

Part A: Write the rate law for the reaction.
A) Rate=k
B) Rate=k[AB]
C) Rate=k[AB]2
D) Rate=k[AB]3

Part B: What is the half-life when the initial concentration is 0.52 M ?
Express your answer using two significant figures.

t1/2 = ?s

Part C: If the initial concentration of AB is 0.230 M , and the reaction mixture initially contains no products, what are the concentrations of A and B after 75 s ?
Express your answers numerically using two significant figures, separated by a comma.
[A],[B] = ?M
1
Expert's answer
2020-06-11T09:22:00-0400

Part A. C) Rate=k[AB]2

Part B. t1/2 = 1 / (kC0) = 1 / (5.5 × 10−2 × 0.52) = 35 s

Part C. [A],[B] = 1 / (kt) - C0 = 0.012 M